A tobacco dealer mixed 12 lbs of one grade of tobacco with 10 lbs of another grade to obtain a blend worth $54. He then made a second blend worth $61 by mixing 8 lbs of the first grade and 15 lbs of the second grade. Find the price per lb in each grade.
choices:
a:) $2 and $3
a:) $2 and $4
a:) $1 and $2
a:) $1 and $3

Answer by mananth(16946)

A tobacco dealer mixed 12 lbs of one grade of tobacco with 10 lbs of another grade to obtain a blend worth $54. He then made a second blend worth $61 by mixing 8 lbs of the first grade and 15 lbs of the second grade. Find the price per lb in each grade.

let blend one price be $x
Let blend 2 price be y
12x+10y=54
8x+15y=61
2(12x+10y) - 3(8x+15y)= 2*54 - 3*61
24x+20y-24x-45y= -75
-25y=-75
y=3.... $3 for the secon blend
plug the value of y in equation 12x+10y=24
x=2
$2, $3 is the price